pooja195 (pooja195):

A van travels 80 miles on 4 gallons of gas. At that rate, how much gas will it need to travel 200 miles? [A] 10 gal [B] 1.6 gal [C] 9 gal [D] 12 gal

OpenStudy (whpalmer4):

If it travels 80 miles on 4 gallons, how many miles per gallon does it get? Once you know that, divide 200 miles by the miles per gallon figure to find the number of gallons.

OpenStudy (whpalmer4):

Another approach: 80:4 is as 200:x Cross-multiply and solve for x \[\frac{80}{4} = \frac{200}{x}\]\[80x = 200*4\]\[x=?\]

pooja195 (pooja195):

800

OpenStudy (whpalmer4):

Are you sure about that? \[80x = 800\]\[80*800 = 800\]That answer doesn't quite seem to fit.

OpenStudy (whpalmer4):

If \[80x=800\]divide both sides of the equation by 80 \[\frac{80x}{80} = \frac{800}{80}\]Cancel common factors\[x=\frac{800}{80} = ?\]

pooja195 (pooja195):

10 right??

OpenStudy (whpalmer4):

Bingo!

pooja195 (pooja195):

yay thnks